IPP Progress Report for the week 2009.12.14 - 2009.12.20

(Up to IPP Progress Reports)

Eugene Magnier

Heather Flewelling

  • culled raw exposures (~690,000) off of machines.
  • fixed up regrevert script so that we only revert nebulous-type errors on gpc1 and so we revert isp registration data. I'm curious if we get the other errors, and if so, if reverting does anything (I'll revert those by hand- we don't often get them).
  • processed 45 exposures to camera stage for Tonry.
  • processing MD03 and SVS with new dark for camera group
  • deleted most of the megacam files off of the cluster (once verified they are in nebulous)
  • worked on processing tasks wiki - it's up to date as far as what's been released (that I know of, unless there are labels I haven't processed). Still thinking about how to better organize it.
  • rebuilt ipp and ipp-debug on ipp-user.
  • trip to maui on friday to see summit operations

Bill Giebink

Roy Henderson

  • Continuing work on IPP->PSPS interface
    • FrameMeta table now written out
    • more columns written to Detection and ImageMeta tables
    • linking with Bill's DVO utility functions, grabbing object/detection IDs etc
    • supplied PSPS with a demo file
    • short week due to 'meeting' on Maui on Friday

Paul Price

Bill Sweeney

  • finished integrating and testing the dvo util library for ipp-psps tool
  • added code to generate statistics during magic destreaking. updated db schema
  • worked on ipp tools enhancements to support postage stamp server queuing update runs for cleaned data
  • analyzed magic streak detection results for data taken after recent camera changes
  • out sick one day

Chris Waters

  • Summit copy: fixed issue with switching databases that was causing data to be downloaded uncompressed.
  • New Dark: Finished reverting problems that were preventing this from building correctly. Looks like it matches the recent data better than the old dark did.
  • Burntool: example of chip with failures has been forwarded to JT for further study.
  • Nebulous: Switched nebulous database from ippdb02 to ippdb00.
  • Cosmic Ray masking: Still not complete, working out bugs in the algorithm.